OpenStudy (anonymous):

need answer 7x+x(x-4)=0

OpenStudy (anonymous):

what are the solutions

OpenStudy (anonymous):

you can see just by looking at it that x=0 is a solution than you can divide by x to get 7+x-4=0 x=-3 so solutions 0 and -3

OpenStudy (anonymous):

maybe clearer if you factor it: x(x+7-4)=x(x+3)=0

OpenStudy (anonymous):

questions says to solve it what are the soultions

OpenStudy (anonymous):

x=-3 and x = 0. This can be found using the solve function on the calculator.

OpenStudy (anonymous):

This is what Andras what saying indirectly. b/c if you move the 3 to the right side of the = you have x = -3 and you also have x = 0.